OpenStudy (anonymous):

Find S7 for the geometric series 2 + -6 + 18 + -54 +… Answer 162 -486 1094 -4374

OpenStudy (cwrw238):

S7 = a1 * r^6 where a1 = first term and r = common ratio

OpenStudy (cwrw238):

r = second term / first term

OpenStudy (anonymous):

So 2*-3^6?

OpenStudy (cwrw238):

oops sorry - i didnt read it properly- you need the sum of 7 terms not the 7th term

OpenStudy (cwrw238):

the formula for S7 is a1 * r^7 - 1 ------- r - 1

OpenStudy (anonymous):

Gimme a second...

OpenStudy (anonymous):

The answer is 1094 :)

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (cwrw238):

its worth commiting that formula to memory sum of n terms Sn = a1 * r^n - 1 ------- r - 1